35. In a distillation set up a) Cold water flows into the condenser from the elevated end of the condenser and flows out of the lower end b) Plastic joint clips should not be used to hold the glass joints of the equipment together, equipment clamps are sufficient c) In a properly set up distillation experiment azeotropes will never occur. d) A and C alone e) None of the above
34. If a distillation is carried out under reduced pressure which of the following point(s) is/are valid: can be obtained at lower distillation temperatures a) Compounds with high boiling points b) The most volatile compounds will be distilled last c) Boiling stones/chips d) Both (a) and (b) e) None of the above are not necessary
39. What is the appropriate way to add the concentrated H3P04 to the alcohol solution to initiate the dehydration reaction? Place the alcohol in a round bottom flask and add the acid dropwise at room temperature. Place the acid in a round bottom flask and add the alcohol dropwise at room temperature Mix the alcohol and acid directly in the round bottom flask. Dropwise addition is too slow. Place the alcohol in the round bottom flask, heat to boiling, then add the acid dropwise Pre-heat the concentrated acid then add to the boiling alcohol directly a) b) d)
18. Which recrystallization solvents would be best for the following compounds? OH H OH Ho HO HO- OH OH H H 3 2 1 a) 1. ethanol; 2. acetone; 3. Acetone b) 1. water; 2. benzene; 3. Ethanol c) 1. water; 2. acetone; 3. diethyl ether d) . acetone; 2. benzene; 3. Water e) None of these systems will work
14. You are taking from 3500- 3200? an IR of a wet compound. What common stretch will you see а) С-Н b) О-Н с) С-BR d) C C еe) С-О
29. Suppose a compound has a distribution coefficient, K = 2, between two solvents. If there are 300 particles of that compound in solution. How many particles will be extracted in two 100 mL extractions vs. one 200 mL more extraction? a) 30 b) 27 с) 90 d) 54 e) 9
Answer #1

Question 35.

The answer is option B.

Explanation: Plastic joint clips should not be used to hold the glass joints of the equipment together, equipment clamps are sufficient.

Question 34.

The answer is option D.

Explanation: Compounds of high boiling points can be obtained at lower distillation temperatures. The most volatile compounds will be distilled last.

Question 39.

The answer is option A.

Explanation: Place the alcohol in a round bottom flask and add the acid dropwise at room temperature.

Question 18.

The answer is option B.

Explanation: 1 (glucose)-water, 2 (pyrene)-benzene, 3 (1-naphthoic acid)-ethanol

Question 14.

The answer is option B.

Explanation: In the IR spectrum, O-H stretching corresponds to 3500-3200 cm-1.

Question 29.

The answer is option E.

Explanation: Final/initial = 2*200/(2*200 + 200) = 2/3 (200 mL extraction)

Particles extracted = 300 - (300*2/3) = 100

Final/initial = {2*200/(2*200 + 100)}2 = 0.64 (2*100 mL extractions)

Particles extracted = 300 - (300*0.64) = 108

Therefore, the more no. of particles extracted = 108 - 100 = 8 ~ 9